Gerald has ten shirts. He needs to pack four of them for a business trip. How many different combinations of shirts could Gerald pack?
Answer by Alan3354(69443)

The 1st selection is 1 of 10, then 1/9, then 1/8, etc
--> 10*9*8*7 = 5040
-----------
But, choosing A,B,C & D is the same as B,D,A & C
There are 4*3*2*1 = 24 ways to get each group.
5040/24 = 210 combos